# Random Matrix Theory-guided sparse PCA for single-cell RNA-seq data

### v1

@@ IV Materials and Methods
$g_{2}(z)$ $=\int\frac{t\rho_{A}(t)}{-z(1+tg_{1}(z))}dt,\quad\forall z\in\mathbb{C}^{+}.$ (13)
@@
Mapping for left-whitened data
For left-whitened data $X\leftarrow A^{-1/2}X$, any outlier eigenvalue $\lambda\notin\operatorname{supp}\rho_{S}$ corresponds to a signal eigenvalue $\alpha\notin\operatorname{supp}\rho_{B}$ of $\mathbb{E}[S]$, given by
@@
where $\psi$ is the functional inverse of $1/\underline{m}(\lambda)$ (see App. A) [26]. Reciprocally, for a signal eigenvector $u$ with eigenvalue $\alpha$, the same relation holds for $\|Wu\|^{2}$, where $W$ is the orthogonal projector onto the outlier eigenspace of $S$ [29]. This result also holds with $Q=uu^{T}$ and $W=vv^{T}$, provided their eigenvalues are simple and sufficiently separated from the other eigenvalues [29, 28]. For biwhitened data $\rho_{B}(t)=\delta(t-1)$ and the Stieltjes transform $\underline{m}(\lambda)$ is given by Eq. 15, and $\psi$ is given by:
@@
Datasets
We downloaded seven publicly available datasets spanning seven single-cell RNA-seq technologies: 10X Chromium v1 (Zheng2017) [37], Multiome (Luecken2021) [35], inDrops (Hrvatin2018) [38], Smart-Seq3xpress (Jensen2022) [36], Drop-Seq (Macosko2015) [39], CITE-seq [40], and 10X Chromium v3 (Wang2020) [41]. For the Multiome dataset, we used only the RNA-seq modality. The dataset were not preprocessed prior to applying the pipelines described in Fig. S2. The only difference across the main text figures is that $2500$ highly variable genes were used for Fig. 1, while $2000$ highly variable genes were used for all other figures. Complete details about data processing are provided in Section B.
@@ Appendix A Almost sure limits of outlier eigenvalues in the separable covariance model
We can rewrite the system of equation for $g_{1}$ and $g_{2}$ shown in Eq. 13 in the form of a single equation, with $g_{2}(z)$ the unique solution $g_{2}$ of $F(z,g_{2})=0$ where $F$ reads
@@
With the knowledge of $\rho_{A}$ and $\rho_{B}$, assuming we can exactly solve this equation, we have access to $m(z)$ via Eq. 11, and by the inversion formula to $\rho_{S}$ and its support.
@@ A.1 Spiked separable covariance model
This equation for $\lambda$ is to be solved outside of $\mathrm{supp}\rho_{S}$, and if there is no solution then it means that $\lambda$ falls on one of the edges $\rho_{S}$. We note that this result has only be proven for eigenvalues $\lambda$ above the rightmost edge of the spectrum [28], but numerical investigation (see Fig. S1) indicate that this result is more general. When there is a solution, there is no guarantee as to whether it is unique, and this equation can have multiple solutions, as also shown in Fig. S1. When $A$ is a low-rank deformation of the identity, i.e. $\rho_{A}(t)=\delta(t-1)$, S2 reduces to the classical result for spiked covariance matrices:
@@ A.2 Information-plus-noise model with separable covariance
Figure S1: Outlier eigenvalues for mixtures of information-plus-noise and separable spiked covariance models. In this figure we use $n=3180$ and $p=3990$. a, b. We define a vector $u_{1}\in\mathbb{R}^{p}$ with $u_{1,1}\approx 0.24$, $u_{1,2}\approx 0.97$, and all other entries zero. For the correlated mixture, we define a vector $u_{2}\in\mathbb{R}^{p}$ with $u_{2,1}\approx 0.92$, $u_{2,2}\approx 0.39$, and all other entries zero, while for independent mixtures $u_{2}$ has entries chosen uniformly at random. We also define a vector $u_{3}\in\mathbb{R}^{n}$ with entries chosen uniformly at random. All vectors are normalized. The matrices $A$ and $B$ are diagonal. We take $B$ with $60\%$ of its entries equal to $12$ and the rest equal to $1$. In a, b, $30\%$ of the entries of $A$ are set to $8$ and the rest to $0.1$. In c, d, AA is the identity matrix. The low-rank signals are defined as $Q=45u_{1}u_{1}^{T}$ and $P=2.5u_{3}u_{2}^{T}$. The data is generated as $X=A^{1/2}Y(B+Q)^{1/2}+P$ with entries of $Y$ being independent standard normal variables. The same realization $Y$ is used across all four subfigures. a, c. Independent mixture. b, d. Correlated mixture.
@@ Appendix B Reproducibility
Each horizontal branch from the main (left-side) pipeline corresponds to an independent copy of the data. Each copy is then processed through the different downstream pipelines, corresponding to the methods shown in Fig. 3 and Fig. S10S13. This design ensures that all methods receive exactly the same data, with identical features and cells. The red arrows indicate that the set of genes selected at earlier steps (quality control and highly variable gene filters) is reused across pipelines. This ensures that the full dataset is processed using the same set of genes as the subsampled dataset. Otherwise, the set of genes selected on the full dataset would differ substantially from that obtained on the subsample. These precautions are necessary to guarantee that each method is evaluated fairly and not biased by differences in gene selection.
The final step of each pipeline is a low-dimensional embedding of the cells. For scVI and DCA, each dataset produces two distinct low-dimensional embeddings (PCA and latent) from the same data copy, though the latent embedding is not shown in the flow chart. The parameters displayed correspond to one realization of the trials in Fig. 3, and the number of components used in the PCA steps may vary slightly. Since DCA and scVI are hyperoptimized, the parameter set for these methods may differ from one trial to another.
Algorithm 2 FISTA sparse PCA
@@
Figure S12: $k$-NN classification performance with different gene sets. Cell type classification performance evaluated as in Fig. 3 and Fig. S10, but using the parameter flavor=’seurat_v3’ in the scanpy package [53], see Fig. S2.
Figure S13: $k$-NN classification performance with $10000$ highly variable genes. Cell type classification performance evaluated as in Fig. 3 and Fig. S10, but using $n=10000$ highly variable genes.

### v2

@@ 4.1 Eigenspectrum of the sample covariance matrix
$g_{2}(z)$ $=\int\frac{t\rho_{A}(t)}{-z(1+tg_{1}(z))}dt,\qquad\forall z\in\mathbb{C}^{+}.$ (13)
@@
4.2 Mapping for left-whitened data
For left-whitened data $X_{\mathrm{lw}}=A^{-1/2}X$, any outlier eigenvalue $\lambda\notin\operatorname{supp}\rho_{S}$ corresponds to a signal eigenvalue $\alpha\notin\operatorname{supp}\rho_{B}$ of $\mathbb{E}[S]$, given by
@@ 4.2 Mapping for left-whitened data
where $\psi$ is the functional inverse of $1/\underline{m}(\lambda)$ (see SI Appendix A) [51]. Reciprocally, for a signal eigenvector $u$ with eigenvalue $\alpha$, the same relation holds for $\|Wu\|^{2}$, where $W$ is the orthogonal projector onto the outlier eigenspace of $S$ [9]. This result also holds with $Q=uu^{T}$ and $W=vv^{T}$, provided their eigenvalues are simple and sufficiently separated from the other eigenvalues [9, 16]. For biwhitened data $\rho_{B}(t)=\delta(t-1)$ and the Stieltjes transform $\underline{m}(\lambda)$ is given by Eq. 15, and $\psi$ is given by:
@@
4.3 Datasets and reproducibility
We use seven publicly available datasets spanning seven single-cell RNA-seq technologies: 10X Chromium v1 (Zheng2017) [53], Multiome (Luecken2021) [36], inDrops (Hrvatin2018) [21], Smart-Seq3xpress (Jensen2022) [20], Drop-Seq (Macosko2015) [38], CITE-seq (Stuart2019) [46], and 10X Chromium v3 (Wang2020) [49]. For the Multiome dataset, we used only the RNA-seq modality. The dataset were not preprocessed prior to applying the pipelines described in Fig. S2. The only difference across the main text figures is that $2500$ highly variable genes were used for Fig. 1, while $2000$ highly variable genes were used for all other figures. Complete details about data processing are provided in SI Appendix B. The code necessary to reproduce the figures of this paper, as well as python implementations of the Alg. 1 and Alg. S1, are available in the following github repository: https://github.com/vchz/spcarmt. For both algorithms we use a threshold over the relative improvement on objective functions as stopping criteria.
5 Acknowledgements
We thank Michael Shelley, the Biophysical Modeling Group, and the Genomics Group at the Flatiron Institute for valuable discussions. We are also grateful to Giulia Pisegna for feedback on the manuscript. The Flatiron Institute is a division of the Simons Foundation.
@@ Appendix A Almost sure limits of outlier eigenvalues in the separable covariance model
We can rewrite the system of equation for $g_{1}$ and $g_{2}$ in the form of a single equation, with $g_{2}(z)$ the unique solution $g_{2}$ of $F(z,g_{2})=0$ where $F$ reads
@@
With the knowledge of $\rho_{A}$ and $\rho_{B}$, assuming we can exactly solve this equation, we have access to $m(z)$, and by the inversion formula to $\rho_{S}$ and its support.
@@ A.1 Spiked separable covariance model
This equation for $\lambda$ is to be solved outside of $\mathrm{supp}\rho_{S}$, and if there is no solution then it means that $\lambda$ falls on one of the edges $\rho_{S}$. We note that this result has only be proven for eigenvalues $\lambda$ above the rightmost edge of the spectrum [16], but numerical investigation (see Fig. S1) indicate that this result is more general. When there is a solution, there is no guarantee as to whether it is unique, and this equation can have multiple solutions, as also shown in Fig. S1. When $A$ is a low-rank deformation of the identity, i.e. $\rho_{A}(t)=\delta(t-1)$, Eq. S2 reduces to the classical result for spiked covariance matrices:
@@ A.2 Information-plus-noise model with separable covariance
Figure S1: Outlier eigenvalues for mixtures of information-plus-noise and separable spiked covariance models. In this figure we use $n=3180$ and $p=3990$. We define a vector $u_{1}\in\mathbb{R}^{p}$ with $u_{1,1}\approx 0.24$, $u_{1,2}\approx 0.97$, and all other entries zero. For the correlated mixture, we define a vector $u_{2}\in\mathbb{R}^{p}$ with $u_{2,1}\approx 0.92$, $u_{2,2}\approx 0.39$, and all other entries zero, while for independent mixtures $u_{2}$ has entries chosen uniformly at random. We also define a vector $u_{3}\in\mathbb{R}^{n}$ with entries chosen uniformly at random. All vectors are normalized. The matrices $A$ and $B$ are diagonal. We take $B$ with $60\%$ of its entries equal to $12$ and the rest equal to $1$. The low-rank signals are defined as $Q=45u_{1}u_{1}^{T}$ and $P=2.5u_{3}u_{2}^{T}$. The data is generated as $X=A^{1/2}Y(B+Q)^{1/2}+P$ with entries of $Y$ being independent standard normal variables. The same realization $Y$ is used across all four subfigures. a, b, $30\%$ of the entries of $A$ are set to $8$ and the rest to $0.1$. The signal eigenvalues $\alpha$ are computed as the isolated eigenvalues of $(B+Q)$. c, d, AA is the identity matrix. The signal eigenvalues $\alpha$ are computed as the isolated eigenvalues of $\mathbb{E}[S]=I+Q+P^{T}P/n$. a, c. Independent mixture. b, d. Correlated mixture.
@@ Appendix B Reproducibility
Each horizontal branch from the main (left-side) pipeline corresponds to an independent copy of the data. Each copy is then processed through the different downstream pipelines. This design ensures that all methods receive exactly the same data, with identical features and cells. The red arrows indicate that the set of genes selected at earlier steps (quality control and highly variable gene filters) is reused across pipelines. This ensures that the full dataset is processed using the same set of genes as the subsampled dataset. Otherwise, the set of genes selected on the full dataset would differ substantially from that obtained on the subsample. These precautions are necessary to guarantee that each method is evaluated fairly and not biased by differences in gene selection.
The final step of each pipeline is a low-dimensional embedding of the cells. For scVI and DCA, each dataset produces two distinct low-dimensional embeddings (PCA and latent) from the same data copy, though the latent embedding is not shown in the flow chart. The parameters displayed correspond to one realization of the trials and the number of components used in the PCA steps may vary slightly. Since DCA and scVI are hyperoptimized, the parameter set for these methods may differ from one trial to another.
Algorithm S1 FISTA sparse PCA
@@
Figure S12: $k$-NN classification performance with different gene sets. Cell type classification performance evaluated using the parameter flavor=’seurat_v3’ in the scanpy package [50], see Fig. S2.
Figure S13: $k$-NN classification performance with $10000$ highly variable genes. Cell type classification performance evaluated Fig. S10, but using $n=10000$ highly variable genes.